# Perturbation-based Effect Measures for Compositional Data

### v2

@@ 4.2.1 Effects of diversity on income in semi-synthetic US census data
We then sample groups from each category, using a weighted sampling approach described in detail in Section S4.3 of the supplementary material. This results in $978$ observations with $811$ in category 00, $92$ in category $1$ and $75$ in category $2$ when grouping on education and $782$ in category 00, $129$ in category $1$ and $67$ in category $2$ when grouping on compensation. For each type of grouping the racial composition for the resulting groups is illustrated in the top half of Figure 5. In both cases we expect that – by construction – there is a positive effect of increased diversity on compensation (for the grouping by race and education this is because education is positively associated with compensation).
Figure 5: Top: Racial compositions for both types of grouping in a ternary plot. Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red).
Bottom: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle). Depending on the color the dependence between $Y$ and the Gini coefficient changes.
@@ 4.3 Variable influence measures when predicting BMI from gut microbiome
Next, we empirically demonstrate that simply ignoring the simplex constraint is not an option. To this end we consider a standard approach for estimating effects of coordinates for unconstrained Euclidean data. More specifically, we compare the proposed $\theta_{j}=\mathrm{CFI}_{\mathrm{unit}}^{j}$ (with estimator based on the partially linear model) with a similar estimator but based on the partially linear model $\mathbb{E}[Y\,|\,Z]=\theta_{j}Z^{j}+h(Z^{-j})$ in which the parameter $\theta_{j}$ is unidentified due to the simplex constraint. For both methods we use a regression estimator that uses cross-validation to select the best estimator out of a linear regression, random forest and support vector regression. As expected the results from the estimator that ignores the simplex structure are meaningless; for example, $\mathrm{var}(Z^{j})\widehat{\theta}_{j}^{2}/\mathrm{var}(Y)$, a proxy for the variance explained by the linear effect, is larger than $10^{17}$ for all species. In contrast, for $\mathrm{CFI}_{\mathrm{unit}}$ these values lie between $10^{-7}$ and $1.6$.
@@
Figure 7: Effect estimates for the $10$ most significant (smallest $p$-value) species from $\mathrm{CFI}_{\mathrm{mult}}$ (top left) and a log-contrast model with pseudocount equal to minimum non-zero observation of $Z$ over $2$ (top right). The labels are colored by the size of the $p$-value. The first $5$ species remain below $0.05$ after Bonferroni-correction for $\mathrm{CFI}_{\mathrm{mult}}$ while this is only true for $2$ for the log-contrast model. The bottom left plot shows the $p$-values from different methods for the three most significant species from each method. We see that the log-contrast model is mostly selecting the same species as the $\mathrm{CKE}$ which explicitly captures effects of zeros. The bottom right plot shows the log-contrast estimates scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$ for the species listed on the top right plot as a function of the added pseudocount with the grey dashed line indicating the pseudocount used in the remaining experiments. We see that the relative influences of the different species is highly contingent on the pseudocount.
@@ 6 Discussion
In this work we have focused on data constraints induced by the simplex. There are however many other types of constraints that can arise, e.g., directional data, general manifold-valued data or otherwise constrained Euclidean data. The idea of defining effect measures via perturbations easily extends to those settings. Similarly, perturbations can also be used to define meaningful effects of summary statistics on unconstrained data; something which has received surprisingly little attention in the literature but relates to problems in causal inference (e.g., how to define causal variables). Nevertheless, for both extensions, more work is required to develop reparametrizations that can be used to derive efficient estimators akin to ours using semiparametric theory. We hope our work will spark new developments in this area and help provide practitioners with the statistical tools they need.
@@ Acknowledgements
The estimators based on the partially linear model are still well-behaved even in settings where the model is misspecified. The estimand $\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)]/\mathbb{E}[\mathrm{var}(L\,|\,W)]$ just happens to be different from $\tau$ and $\lambda$ (depending on the setting). We can see this by plotting the distribution of the estimates from the $d=15$ experiment (with $Y$ nonparametric) of Section 4.1 which we do in Figure S3. Even though the estimand is different, it is still related to the original target parameters, since it can be seen as the ’best approximation’ to a generic nonparametric model, as shown in the following proposition.
@@
Figure S3: Smoothed empirical distributions of the estimates for the nonparametric model (NPM) and partially linear model (PLM) estimators in the nonparametric settings described in Section 4.1 with $d=15$. We see that while the NPM estimator correctly clusters around the true value of $1$, the PLM estimator also clusters however around a different, but potentially still useful, value.
@@
Section S4.1 provides the details behind the experiments in Table 1 in Section 1. Section S4.2 contains additional simulation results from our investigation into the performance of various semiparametric estimators. Section S4.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.2.1. Section S4.5 provides a detailed description of the non-compositional background variables that are controlled for in Section 4.2.2.
@@
Let $L\sim\mathrm{Bern}(1/2)$, $U^{W},U^{L}\sim\mathrm{Unif}([0,1])$ and $B\sim\mathrm{Bern}(1/2)$ be independent random variables and define
@@
where $W^{2}:=U^{W}(1-L)+U^{W}LB$ and $Z^{1}:=(1-L)U^{L}$. Furthermore, define $Y$ such that
$Y\,|\,Z\sim\mathrm{Bern}\left(\frac{3}{4}-\frac{1}{8}\mathds{1}_{\{Z^{1}=0\}}+\frac{1}{2}\mathds{1}_{\{Z^{2}=0\}}\right).$
@@
S4.2 Additional semiparametric estimation results
Here we provide additional results from the experiment of Section 4.1. Figure S4 and Figure S5 contain estimated coverage rates of the $95\%$ confidence intervals for the different methods and settings with $d=3$ and $d=75$, respectively. The patterns observed in Figure 4 of Section 4.1 are repeated in these figures.
Figure S4: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section 4.1 with $d=3$. The patterns observed in Figure 4 of Section 4.1 are repeated here.
Figure S5: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section 4.1 with $d=75$. The patterns observed in Figure 4 of Section 4.1 are repeated here.
@@
S4.4 Semi-synthetic individual-level experiment using groups from Section 4.2.1
@@
S4.5 Description of additional control variables used in Section 4.2.2
@@
S4.6 Additional plots from gut microbiome analysis in Section 4.3

### v3

@@ 4.1.1 Effects of diversity on income in semi-synthetic US census data
We then sample groups from each category, using a weighted sampling approach described in detail in Section S4.3 of the supplementary material. This results in $978$ observations with $811$ in category 00, $92$ in category $1$ and $75$ in category $2$ when grouping on education and $782$ in category 00, $129$ in category $1$ and $67$ in category $2$ when grouping on compensation. For the data where we have grouped on education the racial composition for the resulting groups is illustrated in the left plot of Figure 4. We expect that – by construction – there is a positive effect of increased diversity on compensation because education is positively associated with compensation. This can be seen in the right plot of Figure 4 which shows how the overall relationship between diversity and compensation is negative but conditioned on $\frac{Z-z_{\mathrm{cen}}}{\|Z-z_{\mathrm{cen}}\|_{1}}$ the effect becomes positive. The same occurs when grouping on compensation and the corresponding figure can be found in Section S4.4 of the supplementary material.
For each type of grouping the racial composition for the resulting groups is illustrated in the top half of Figure 4. In both cases we expect that – by construction – there is a positive effect of increased diversity on compensation (for the grouping by race and education this is because education is positively associated with compensation). This can be seen in the lower half of Figure 4 which shows how the overall relationship between diversity and compensation is negative but conditioned on $\frac{Z-z_{\mathrm{cen}}}{\|Z-z_{\mathrm{cen}}\|_{1}}$ the effect becomes positive.
Figure 4: Summary of the semi-synthetic ‘Adult’ dataset when grouping on education. Left: Racial compositions in a ternary plot. Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red). Right: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle). Depending on the color the dependence between $Y$ and the Gini coefficient changes. The Gini coefficient represents a similarity rather than a diversity measure.
@@ 4.2 Variable influence measures when predicting BMI from gut microbiome
Next, we empirically demonstrate that simply ignoring the simplex constraint is not an option. To this end we consider a standard approach for estimating effects of coordinates for unconstrained data. More specifically, we compare the proposed $\theta_{j}=\mathrm{CFI}_{\mathrm{unit}}^{j}$ (with estimator based on the partially linear model) with a similar estimator but based on the partially linear model $\mathbb{E}[Y\,|\,Z]=\theta_{j}Z^{j}+h(Z^{-j})$ in which the parameter $\theta_{j}$ is unidentified due to the simplex constraint. For both methods we use a regression estimator that uses cross-validation to select the best estimator out of a linear regression, random forest and support vector regression. As expected the results from the estimator that ignores the simplex structure are meaningless; for example, $\mathrm{var}(Z^{j})\widehat{\theta}_{j}^{2}/\mathrm{var}(Y)$, a proxy for the variance explained by the linear effect, is larger than $10^{17}$ for all species. In contrast, for $\mathrm{CFI}_{\mathrm{unit}}$ these values lie between $10^{-7}$ and $1.6$.
@@
Figure 6: Effect estimates for the $10$ most significant (smallest $p$-value) species from $\mathrm{CFI}_{\mathrm{mult}}$ (top left) and a log-contrast model with pseudocount equal to minimum non-zero observation of $Z$ over $2$ (top right). The labels are colored by the size of the $p$-value. The first $5$ species remain below $0.05$ after Bonferroni-correction for $\mathrm{CFI}_{\mathrm{mult}}$ while this is only true for $2$ for the log-contrast model. The bottom left plot shows the $p$-values from different methods for the three most significant species from each method (text color only changes to enhance readability). We see that the log-contrast model is mostly selecting the same species as the $\mathrm{CKE}$ which explicitly captures effects of zeros. The bottom right plot shows the log-contrast estimates scaled by the standard deviation of $\log(Z^{j}+\text{pseudocount})$ for the species listed on the top right plot as a function of the added pseudocount with the grey dashed line indicating the pseudocount used in the remaining experiments. We see that the relative influences of the different species is highly contingent on the pseudocount.
@@ 6 Discussion
In this work we have focused on data constraints induced by the simplex. There are however many other types of constraints that can arise, e.g., directional data, general manifold-valued data or otherwise constrained data. The idea of defining effect measures via perturbations easily extends to those settings. Similarly, perturbations can also be used to define meaningful effects of summary statistics on unconstrained data; something which has received surprisingly little attention in the literature but relates to problems in causal inference (e.g., how to define causal variables). Furthermore, there may be settings where the perturbation of interest depends on additional covariates which then changes the resulting target of inference. Nevertheless, for all of the above extensions, more work is required to develop reparametrizations that can be used to derive efficient estimators akin to ours using semiparametric theory. We hope our work will spark new developments in this area and help provide practitioners with the statistical tools they need.
@@ Acknowledgements
The estimators based on the partially linear model are still well-behaved even in settings where the model is misspecified. The estimand $\mathbb{E}[\mathrm{cov}(Y,L\,|\,W)]/\mathbb{E}[\mathrm{var}(L\,|\,W)]$ just happens to be different from $\tau$ and $\lambda$ (depending on the setting). We can see this by plotting the distribution of the estimates from the $d=15$ experiment (with $Y$ nonparametric) of Section S4.2 which we do in Figure S3. Even though the estimand is different, it is still related to the original target parameters, since it can be seen as the ’best approximation’ to a generic nonparametric model, as shown in the following proposition.
@@
Figure S3: Smoothed empirical distributions of the estimates for the nonparametric model (NPM) and partially linear model (PLM) estimators in the nonparametric settings described in Section S4.2 with $d=15$. We see that while the NPM estimator correctly clusters around the true value of $1$, the PLM estimator also clusters however around a different, but potentially still useful, value.
@@
Section S4.1 provides the details behind the experiments in Table 1 in Section 1. Section S4.2 contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional. Section S4.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.1.1. Section S4.4 includes an additional figure showing the semi-synthetic dataset used in Section 4.1.1. Section S4.6 provides a detailed description of the non-compositional background variables that are controlled for in Section 4.1.2.
@@
Let $L\sim\mathrm{Bern}(1/2)$, $U^{1},U^{2}\sim\mathrm{Unif}([0,1])$ and $B\sim\mathrm{Bern}(1/2)$ be independent random variables and define
@@
where $W^{2}:=U^{1}(1-L)+U^{1}LB$ and $Z^{1}:=(1-L)U^{2}$. Furthermore, define $Y$ such that $Y$ is Bernoulli distributed with
$\mathbb{P}(Y=1\,|\,Z)=\sim\mathrm{Bern}\biggl{(}\frac{3}{4}-\frac{1}{8}\mathds{1}_{\{Z^{1}=0\}}+\frac{1}{2}\mathds{1}_{\{Z^{2}=0\}}\biggr{)}.$
@@
S4.2 Semiparametric estimators
  • NPM_no_x: The same estimators as NPM but without cross-fitting.
  • PLM_no_x: The same estimators as PLM but without cross-fitting.
The variance of the NPM and PLM estimates is computed based on the provided algorithms for the NPM and PLM methods while the variance for the plug-in is estimated using an estimator akin to the one given in (13). Each estimator above is itself based on multiple of the regression estimates $\widehat{f}$, $\widehat{\pi}$, $\widehat{\rho}$, $\widehat{g}$ and $\widehat{m}$. For the regression estimates, we use a $5$-fold cross-validation based estimator which for each fit selects among the following three models: a simple model that predicts the mean of the outcome (or the proportion of each class for binary classification), a random forest with $250$ trees with trees grown to a maximum depth of $2$ and a random forest with $250$ trees and no depth restriction. For the score estimator we use the location-scale based estimator described in Section S3.4, that uses the same $5$-fold cross-validated regression estimator for the location and scale nuisance regressions.
  • Binary $L$ with partially linear $Y$:
  • Binary $L$ with nonparametric $Y$:
  • Continuous $L$ with partially linear $Y$:
    $L=B(W_{1})+\xi\quad\text{and}\quad Y=L+B(W_{1})+\varepsilon$
  • Continuous $L$ with nonparametric $Y$:
The noise variables $U_{0}\sim\mathrm{Bern}(0.8)$, $U_{1}\sim\mathrm{Bern}(0.5)$, $\varepsilon\sim\mathcal{N}(0,1)$ and $\xi\sim\mathcal{N}(0,1)$ are all sampled independently (and independent of $W$). The true value of $\lambda$ in the binary $L$ settings and $\tau$ in the continuous $L$ settings is $1$. For each setting we simulate data and compute an estimate with a corresponding asymptotic $95\%$ confidence interval using each of the methods described above. We repeat the experiment $100$ times and compute coverage rates of the $95\%$ confidence intervals for $d=3$ in Figure S4, for $d=15$ in Figure S5 and for $d=75$ in Figure S6.
Figure S4: Estimated coverage rates of the $95\%$ confidence intervals for each different estimation procedure over $100$ repetitions for each of the settings described in Section S4.2 with $d=3$. PLM works well if $Y$ follows a partially linear model, but can fail if this assumption is not satisfied as in the nonparametric $Y$ setting. NPM works in both settings but can be less robust for small sample sizes. Further, it can be seen that the plug-in estimator and the estimators without cross-fitting do not lead to nominal coverage in some of these settings.
Figure S5: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section S4.2 with $d=15$. The patterns observed in Figure S4 are repeated here.
Figure S6: Estimated coverage rates of the $95\%$ confidence intervals over $100$ repetitions for each of the settings described in Section S4.2 with $d=75$. The patterns observed in Figure S4 are repeated here.
@@
S4.4 Additional figure for the ‘Adult’ data experiment in Section 4.1.1
Figure S7 complements Figure 4 in Section S4.4 by summarizing the semi-synthetic ‘Adult’ dataset when grouping on compensation.
Figure S7: Summary of the semi-synthetic ‘Adult’ dataset when grouping on compensation. Left: Racial compositions in a ternary plot. Each point corresponds to one group and is colored according to which category it belongs to (00 in blue, $1$ in green and $2$ in red). Right: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle). Depending on the color the dependence between $Y$ and the Gini coefficient changes. Note that the Gini coefficient represents a similarity rather than a diversity measure.
S4.5 Semi-synthetic individual-level experiment using groups from Section 4.1.1
@@
S4.6 Description of additional control variables used in Section 4.1.2
@@
S4.7 Additional plots from gut microbiome analysis in Section 4.2